whats the difference between division problem,fraction,ratios,percentages?are they all the same things?
Answer by mark_henrhiz(40)

YES. THEY HAVE THE SAME VALUES.
FOR EXAMPLE
1 DIVIDED BY 2 = 0.5 = 5/10 = 1/2 = 50%